A sphere thank of 1.4 m radius is half filled with oil of relative density 0.8. If the tank is given a horizontal acceleration of 10mps2,  the maximum pressure on the tank is root2 pascal. Find the value of p.

Grade:9

1 Answers

Arun
5 years ago
The liquid rises along one of the sides.
so,
Let the rise be h.
So,
h/2r=a/g
=> h= a/g * 2r
Since a=10m/s2
h=2r=2.4m
Therefore maximum pressure on the tank = ρgh
= 800 x 10 x 2.4
= 1.92 x 104 Pa